OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution xapproachesinfinity

OpenStudy (xapproachesinfinity):

hey the same thing here! go back to our problem earlier and see we have done

OpenStudy (anonymous):

how do you start divide both sides by -3?

OpenStudy (anonymous):

No. Distribute the 3.

OpenStudy (anonymous):

The equation would look like -3(2x+6)=-12

OpenStudy (anonymous):

And the other would be 3(2x+6)=-12

OpenStudy (anonymous):

xapproachesinfinity please explain

OpenStudy (xapproachesinfinity):

Faith you are wrong

OpenStudy (anonymous):

How?

OpenStudy (xapproachesinfinity):

@lanM123 you right you divide by -3

OpenStudy (anonymous):

im almost positive that you have to divide both sides by -3

OpenStudy (anonymous):

Why would you divide by -3?

OpenStudy (xapproachesinfinity):

no you are positive no only almost positive ^_^

OpenStudy (anonymous):

so you get |2x+6|=4 correct?

OpenStudy (anonymous):

I got the second answer.

OpenStudy (xapproachesinfinity):

because yo have an absolute value you need to isolate it first otherwise you can't go further those are not parenthesis to do what you said

OpenStudy (xapproachesinfinity):

Yes! correct!

OpenStudy (xapproachesinfinity):

you are going in good direction

OpenStudy (anonymous):

okay what next?

OpenStudy (xapproachesinfinity):

now the critical step!

OpenStudy (xapproachesinfinity):

2x+6=4 or 2x+6=-4

OpenStudy (xapproachesinfinity):

recall what we did! we said that inside the abs value can positive or nagative so if is positive we just say 2x+6=4 if it is negative we have to make it 2x+6=-4 so we are making up for the negative since whatever you through in abs value it will give us a positive resutl it called abs value because of this (aka: the distance! distance can never be negative it will make no since to say i went -3 three step to my left)

OpenStudy (xapproachesinfinity):

sense*

OpenStudy (xapproachesinfinity):

sorry for the grammar

OpenStudy (anonymous):

its okay and this is the part that i don't get

OpenStudy (xapproachesinfinity):

i see you are still having trouble with this! |x|=a ===> x=a or -x=a

OpenStudy (xapproachesinfinity):

no problem with this?

OpenStudy (anonymous):

no problem that is the 2 option thing

OpenStudy (anonymous):